Calculate the amount of heat needed (in kJ with two significant figures) to raise the temperature of 203 g of n-Pentane from 3.37 ° C to 90.00 ° C.
Answer #1

Em = is sol mass of n-pentane 203g = 0-203 kg (as 1kg = 1000 g) Temperature is raised from initial tenperature = Tg = 3-37°C